# Numerical Twin with Two Dimensional Ornstein–Uhlenbeck Processes of Transient Oscillations in EEG signal

### v1

@@ I Introduction
Noise-driven stochastic oscillations arise across physical, chemical, and biological systems. In many settings the observable signal is not a sustained sinusoid but a sequence of transient bursts that recur at irregular intervals and depart markedly from deterministic oscillators. Such behavior has been reported in seismology, biochemistry (e.g., calcium signaling dynamics [Sneyd2017Dynamical]), cardiac electrophysiology [Glass2020], and neural activity, both intracellularly [Gray1996Chattering, Hughes2004Synchronized, Fuentealba2005Membrane, lHorincz2009Temporal] and at the population level via local field potentials (LFP) [buzsaki2006rhythms, adesnik2018cell, adamantidis2019oscillating, veit2023cortical]. In clinical electroencephalography (EEG) [da1973organization], these bursts appear as spindle-shaped events within specific frequency bands; for example, $\delta$ and $\alpha$ spindles are characteristic of sleep [da1973organization, lopes1974model], while fast ripples (100–200 Hz) have been linked to learning and memory consolidation [girardeau2011hippocampal].
Extracting reliable statistics from such nonstationary, narrowband burst patterns remains challenging, particularly for real-time brain monitoring during normal activity, sleep, or general anesthesia (GA) (Fig. 1). Recent CNN/RNN/transformer approaches achieve strong detection on curated sleep datasets, with reported F1 scores typically in the 0.750.750.900.90 range and high score for binary spindle/non-spindle decisions [kulkarni2019spindlenet, muller2022waveform, mofrad2022waveform, Chen2021AutomatedSleepSpindle]. However, these methods generally deliver labels, not physiologically grounded estimators used for prediction, requiring a precise and robust segmentation approach [Chen2021AutomatedSleepSpindle, You2024AnestheticSpindles].
@@
Time–frequency and time-scale analysis (short-time Fourier and wavelet transforms [Mey90I]) quantify spectral content but have limited ability to localize irregular, sporadic oscillatory structures whose envelopes vary slowly in time [jaffard2001wavelets]. Empirical Mode Decomposition (EMD) can better capture modulated components [flandrin2004empirical], yet a compact, generative model that reproduces both event morphology and statistics is still lacking.
We shall use the narrow band two-dimensional Ornstein–Uhlenbeck (OU) process whose linear deterministic part forms a stable focus. While the process is stationary, individual realizations exhibit transient, burst-like oscillations that capture the spindle morphology observed in EEG signals. Noise-driven excursions around this focus generate transient, spindle-like activity [zonca2021emergence]. The model has three interpretable parameters: decay rate $\lambda$, a main frequency $\omega$, and noise amplitude $\sigma$, which together control event rate, coherence, and variability. We shall estimate these parameters from data using two complementary strategies: (i) a global fit that jointly leverages the power spectral density (PSD), marginal amplitude distribution, and autocorrelation function; and (ii) an event-wise fit that segments bursts (via envelope methods related to EMD [huang1998empirical, flandrin2004empirical, jaffard2001wavelets, Wu2009]), then matches empirical spindle statistics (duration, amplitude, and inter-event interval) to OU simulations via grid search, with event counts resolving parameter degeneracies.
@@ II EEG projection on two-dimensional OU-processes
To model burst oscillation, we use a three parameters two-dimensional Ornstein-Uhlenbeck process defined for the state variable $\mbox{$s$}=(x,y)\in\Re^{2}$ as !!!
@@
Our goal is to estimate the three parameters $(\lambda,\omega,\sigma)$ from empirical time series: the optimal estimation constitutes the projection of the signal to the OU-process, with the optimal parameters generating an OU signal close to the EEG one, as shown in Fig.1, for an EEG signal during general anesthesia and sleep and a OU-process realization, which can recapitulate properties of the empirical signal.
@@
is solution of the Fokker-Planck equation (FPE) [schuss2009theory]
@@
While the OU Fokker–Planck operator admits an eigenfunction–eigenvalue expansion for the probability density, such spectral decompositions describe stationary transient distributions and relaxation modes. They do not capture the inherently stochastic, short-lived bursts. Because spindles are defined by localized amplitude excursions bounded by waxing–waning envelopes, their analysis must occur in the joint time–frequency plane. The present inference relies on spectrograms and segmented event statistics rather than on global eigenfunction expansions. The steady-state [risken1996fokker, schuss2009theory] of eq.3 is [risken1996fokker] $f(\mbox{$s$})=\frac{1}{2\pi\sqrt{\det R}}\exp\left(-\frac{1}{2}\mbox{$s$}^{T}R^{-1}\mbox{$s$}\right),$ where the matrix $R$ satisfies Lyapunov’s equation $(-A)R+R(-A)^{T}=2\sigma I_{2},$ where $I_{2}$ is the identity matrix in dimension 2, leading to $R=\frac{\sigma}{\lambda}I_{2}$ and the steady-solution is
@@
At this stage, we note that While the Ornstein–Uhlenbeck process admits well-established asymptotic theories (small-noise WKB expansions [schuss2009], Fokker–Planck eigenfunction decompositions, long-time convergence), these frameworks describe ensemble-averaged behavior and cannot directly extract time-localized transient bursts from individual sample paths. Crucially, as noise amplitude $\sigma\to 0$, the normalized power spectral density (see eq. 10 below) gives
@@
which remains independent of $\sigma$: the resonance structure (peak location $\omega/(2\pi)$ and bandwidth $\lambda$) persists even as the overall amplitude vanishes, ensuring that spindles remain distinguishable from background. Eigenfunction expansions of the Fokker-Planck operator describe ensemble-averaged transient distributions and relaxation modes but miss discrete transient event structure (waxing-waning envelopes, inter-spindle intervals, counts). Moreover small-noise asymptotics [schuss2009] capture rare large deviations rather than the typical resonant fluctuations that generate spindles. We therefore develop here a hybrid approach: 1- the OU theory provides the generative model with explicit parameters $(\lambda,\omega,\sigma)$ constrained by analytic results (Gaussian steady state, autocorrelation, PSD), while the signal processing tools (envelope extraction, segmentation, spectrogram analysis) provide the empirical observables (duration, amplitude, event counts) required to identify and quantify individual bursts. This methodological synthesis is necessary because neither framework alone suffices to bridge stochastic dynamics and time-localized physiological events. Autocorrelation function: The second estimator uses the autocorrelation function of the component $x(t)$ of the OU-process. The power spectral density (PSD) can be computed from the Fourier transform of the autocorrelation function:
@@
Equation (10) shows that the OU focus yields a pair of symmetric Lorentzian peaks centered at $\pm\omega/(2\pi)$ with width $\lambda$. In narrow bands of interest (e.g., $\delta$ and $\alpha$) the empirical PSD is dominated by such Lorentzians plus a slowly varying aperiodic background [loison2024mapping, wen_separating_2016, donoghue2020parameterizing]). Therefore, a multiscale characterization is not required for our goal: we model band-limited, quasi-monochromatic bursts with linewidth and center frequency, rather than broadband cascade dynamics. Though multiscale tools prove useful for cross-band scaling laws [jaffard2001wavelets, dora2024wqn], here we explicitly restrict to narrow band transients, where isolated Lorentzians are sufficient. Substituting the expression for the correlation function (see Appendix) associated to the OU-process
@@ II-A Spindle segmentation and estimating the mean spindle frequency $\hat{\omega}$
To estimate the mean oscillation frequency $\omega$ within each spindle, we first need to segment them. We shall use the following procedure: first, we compute the amplitude envelope of the signal $x(t)$ and then identify discrete spindle events using an empirical thresholding segmentation rule. We start with the classical chirp complex representation $x(t)=A(t)\exp(i\phi(t))$, where $A(t)$ is the instantaneous envelope that varies slowly in time, and $\phi(t)$ is the instantaneous phase capturing the fast oscillatory behavior. The envelope $A(t)$ can computed from the first step of the Empirical Mode Decomposition (EMD) [huang1998empirical, flandrin2004empirical]. We define an isolated spindle pattern as an oscillatory events occurring between two consecutive local minima of the amplitude envelope, separated by a local maximum. Specifically, our segmentation procedure follows the following steps:
@@
    Noise reduction via smoothing: Estimate the upper $x_{\text{up}}(t)$ and lower $x_{\text{down}}(t)$ envelopes by applying locally weighted polynomial regression (LOESS) using a second-degree model for the local maxima and minima, respectively [jacoby2000loess].
@@
Using the segmentation describe above, we extract the spindle duration, peak amplitude, and mean frequency.
@@
where “p.v.” is the Cauchy principal value. From the conjugate component $y_{k}(t)=\mathcal{H}\{x_{k}\}(t)$, we have $x_{a,k}(t)=x_{k}(t)+iy_{k}(t)=A_{k}(t)\exp(i\phi_{k}(t))$, where $A_{k}(t)$ is the instantaneous amplitude and $\phi_{k}(t)$ is the instantaneous phase of spindle $k$. The instantaneous frequency is given by $\omega_{k}(t)=\frac{\mathrm{d}}{\mathrm{d}t}\phi_{k}(t)$, which is a pointwise measure of oscillatory activity within the spindle [daubechies2011synchrosqueezed]. The mean frequency of spindle $k$, segmented between time $t_{k}$ and $t_{k+1}$ is the average over its duration $T_{k}=t_{k+1}-t_{k}$:
@@ III-A General projection on two-dimensional OU-processes
We apply here the OU-decomposition to the EEG signal during general anesthesia, mostly dominated by two main frequency bands: $\delta$ (0.5-4 Hz) and $\alpha$ (8-14 Hz). These bands results from random spindle patterns (Fig. 3A-C).
Figure 3: Spectral decomposition of EEG in $\alpha$- and $\delta$- and residual components A. EEG and spectrogram over several seconds (scal bar 5s), showing $\alpha$- and $\delta$- spindles. B. $\alpha$- and $\delta$- filtered components. C. Power spectral decomposition (PSD). We fitted the PSD in intermediate frequency bands $[1.8,4]$ and [6.5,8.1][6.5,8.1]Hz with $\log PSD=a\log(f)+b$. D. Statistical result of (a,b) coefficients for the two bands and Lorentzian parameters $\frac{A\sigma}{\pi}\frac{1}{\sigma^{2}+(f-f_{c})^{2}}$, computed over n=18 EEGs.
@@
[…]
To conclude we proposed here a decomposition of the EEG signal with OU-processes centered around a dominant frequency band. This approximation are valid on some restricted intervals, separated by scaling law in between.
@@ IV Estimating the OU-parameters based on spindle statistics
Figure 4: Parameter estimation of the stochastic OU- model to EEG spindles via grid search. (A) Parameter estimation pipeline. A grid search is performed over the parameter space of the two-dimensional Ornstein–Uhlenbeck process ($\lambda$, $\sigma$), while fixing the resonance frequency $f_{c}=\omega/2\pi$. (B) Grid search minimizing a loss function based on spindle statistics with ( duration, and amplitude) computed from segmented EEG recordings, this leads to a line of minimum with slope $\lambda/\sigma=0.082$. However, when considering the spindle statistics with (number, duration, and amplitude), the optimal parameters are $\lambda^{*}=1.84$, $\sigma^{*}=35.59$, and $f_{c}^{*}=9.63$ Hz. (C) EEG signal (top) and stochastic simulations of the OU process (bottom) using the optimal parameters. (D) Spindle segmentation applied to both real EEG and simulated signals. Events are identified and classified based on amplitude and duration thresholds. (E) Comparison of spindle statistics (number, amplitude, and duration) between EEG and simulations.
Using a square grid search, For a fixed ensemble $(\lambda,\sigma)$ with predefined boundary, the method consists in generating OU-realizations, segment their spindles to obtain the duration and amplitude distribution and compare them with the distribution of the empirical signal (Fig. 4A). The next point is obtained by $(\lambda+\delta\lambda,\sigma+\delta\sigma)$. To compare the two distribution with the empirical one, we use the Wasserstein distance [friedman2001] based on the sum of absolute differences between the two cumulative density functions
@@
The optimization procedure to implement that allow identify the optimal parameters is
@@
The results of such a procedure (Fig. 4B) depends on the parameter to be considered: when considering the duration and amplitude only in eq. 26, we do not find any single optimal parameter, due to a degeneracy (duration is proportional to amplitude as shown in Fig. 4B-Last raw. In that case, we obtain a line of minimum where the ratio $\frac{\sigma}{\lambda}$ is constant (fig. 4B). However, when adding the number of spindles to the two other parameters, we remove the degeneracy and obtain a global minimum for the loss function $S_{\lambda,\sigma}$, associated with the optimal parameters (fig. 4B). Finally, we obtain a similar trends between the amplitude and duration distribution of the EEG signal vs the optimal reconstructed OU-processes. To conclude such second approach allow to estimate OU-parameters based on the statistics of spindles.

### v2

@@ I Introduction
Noise-driven stochastic oscillations arise across physical, chemical, and biological systems. In many settings the observable signal is not a sustained sinusoid but a sequence of transient bursts that recur at irregular intervals and depart markedly from deterministic oscillators. Such behavior has been reported in seismology, biochemistry (e.g., calcium signaling dynamics [1]), cardiac electrophysiology [2], and neural activity, both intracellularly [3, 4, 5, 6] and at the population level via local field potentials (LFP) [7, 8, 9, 10]. In clinical electroencephalography (EEG) [11], these bursts appear as spindle-shaped events within specific frequency bands; for example, $\delta$ and $\alpha$ spindles are characteristic of sleep [11, 12], while fast ripples (100–200 Hz) have been linked to learning and memory consolidation [13].
Extracting reliable statistics from such nonstationary, narrowband burst patterns remains challenging, particularly for real-time brain monitoring during normal activity, sleep, or general anesthesia (GA) (Fig. 1). Recent CNN/RNN/transformer approaches achieve strong detection on curated sleep datasets, with reported F1 scores typically in the 0.750.750.900.90 range and high score for binary spindle/non-spindle decisions [14, 15, 16, 17]. However, these methods generally deliver labels, not physiologically grounded estimators used for prediction, requiring a precise and robust segmentation approach [17, 18].
@@
Time–frequency and time-scale analysis (short-time Fourier and wavelet transforms [19]) quantify spectral content but have limited ability to localize irregular, sporadic oscillatory structures whose envelopes vary slowly in time [20]. Empirical Mode Decomposition (EMD) can better capture modulated components [21], yet a compact, generative model that reproduces both event morphology and statistics is still lacking.
We shall use the narrow band two-dimensional Ornstein–Uhlenbeck (OU) process whose linear deterministic part forms a stable focus. While the process is stationary, individual realizations exhibit transient, burst-like oscillations that capture the spindle morphology observed in EEG signals. Noise-driven excursions around this focus generate transient, spindle-like activity [22]. The model has three interpretable parameters: decay rate $\lambda$, a main frequency $\omega$, and noise amplitude $\sigma$, which together control event rate, coherence, and variability. We shall estimate these parameters from data using two complementary strategies: (i) a global fit that jointly leverages the power spectral density (PSD), marginal amplitude distribution, and autocorrelation function; and (ii) an event-wise fit that segments bursts (via envelope methods related to EMD [23, 21, 20, 24]), then matches empirical spindle statistics (duration, amplitude, and inter-event interval) to OU simulations via grid search, with event counts resolving parameter degeneracies. Beyond classical second-order descriptor, we will show that spindle statistics (counts, durations, and amplitude–duration structure) impose constraints that are not captured by second-order structure alone, and are essential to resolve parameter in transient oscillatory bursts.
@@ II EEG projection on two-dimensional OU-processes
To model burst oscillation, we use a three parameters two-dimensional Ornstein-Uhlenbeck process defined for the state variable $\mbox{$s$}=(x,y)\in\Re^{2}$ as
@@
Note that Because eq. 1 is a linear Gaussian SDE, the sampled process admits an exact discrete-time representation. Thu, an exact Gaussian likelihood for jointly estimating $(\lambda,\omega,\sigma)$ is available, but in this work segmentation is used to extract event-level spindle statistics beyond second-order structure. we will show that event-level spindle statistics (counts, durations, and amplitude–duration structure) impose constraints that are not captured by second-order structure alone, and are essential to resolve parameter in transient oscillatory bursts. Thus our goal is to estimate the three parameters $(\lambda,\omega,\sigma)$ from empirical time series in the context of segmented spindle events: the optimal estimation constitutes the projection of the signal to the OU-process, with the optimal parameters generating an OU signal close to the EEG one, as shown in Fig.1, for an EEG signal during general anesthesia and sleep and a OU-process realization, which can recapitulate properties of the empirical signal.
@@
is solution of the Fokker-Planck equation (FPE) [28]
@@
While the OU Fokker–Planck operator admits an eigenfunction–eigenvalue expansion for the probability density, such spectral decompositions describe stationary transient distributions and relaxation modes. They do not capture the inherently stochastic, short-lived bursts. Because spindles are defined by localized amplitude excursions bounded by waxing–waning envelopes, their analysis must occur in the joint time–frequency plane. The present inference relies on spectrograms and segmented event statistics rather than on global eigenfunction expansions. The steady-state [29, 28] of eq.3 is [29] $f(\mbox{$s$})=\frac{1}{2\pi\sqrt{\det R}}\exp\left(-\frac{1}{2}\mbox{$s$}^{T}R^{-1}\mbox{$s$}\right),$ where the matrix $R$ satisfies Lyapunov’s equation $(-A)R+R(-A)^{T}=2\sigma I_{2},$ where $I_{2}$ is the identity matrix in dimension 2, leading to $R=\frac{\sigma}{\lambda}I_{2}$ and the steady-solution is
@@
At this stage, we note that, while the Ornstein–Uhlenbeck process admits well-established asymptotic theories (small-noise WKB expansions [30], Fokker–Planck eigenfunction decompositions, long-time convergence), these frameworks describe ensemble-averaged behavior and cannot directly extract time-localized transient bursts from individual sample paths. Crucially, as noise amplitude $\sigma\to 0$, the normalized power spectral density (see eq. 10 below) gives
@@
which remains independent of $\sigma$: the resonance structure (peak location $\omega/(2\pi)$ and bandwidth $\lambda$) persists even as the overall amplitude vanishes, ensuring that spindles remain distinguishable from background. Eigenfunction expansions of the Fokker-Planck operator describe ensemble-averaged transient distributions and relaxation modes but miss discrete transient event structure (waxing-waning envelopes, inter-spindle intervals, counts). Moreover small-noise asymptotics [30] capture rare large deviations rather than the typical resonant fluctuations that generate spindles. We therefore develop here a hybrid approach: 1- the OU theory provides the generative model with explicit parameters $(\lambda,\omega,\sigma)$ constrained by analytic results (Gaussian steady state, autocorrelation, PSD), while the signal processing tools (envelope extraction, segmentation, spectrogram analysis) provide the empirical observables (duration, amplitude, event counts) required to identify and quantify individual bursts. This methodological synthesis is necessary because neither framework alone suffices to bridge stochastic dynamics and time-localized physiological events. Autocorrelation function: The second estimator uses the autocorrelation function of the component $x(t)$ of the OU-process. The power spectral density (PSD) can be computed from the Fourier transform of the autocorrelation function:
@@
Equation (10) shows that the OU focus yields a pair of symmetric Lorentzian peaks centered at $\pm\omega/(2\pi)$ with width $\lambda$. In narrow bands of interest (e.g., $\delta$ and $\alpha$) the empirical PSD is dominated by such Lorentzians plus a slowly varying aperiodic background [31, 32, 33]). Therefore, a multiscale characterization is not required for our goal: we model band-limited, quasi-monochromatic bursts with linewidth and center frequency, rather than broadband cascade dynamics. Though multiscale tools prove useful for cross-band scaling laws [20, 34], here we explicitly restrict to narrow band transients, where isolated Lorentzians are sufficient. Substituting the expression for the correlation function (see Appendix) associated to the OU-process
@@ II-A Spindle segmentation and estimating the mean spindle frequency $\hat{\omega}$
To estimate the mean oscillation frequency $\omega$ within each spindle, we first need to segment them. We shall use the following procedure: first, we compute the amplitude envelope of the signal $x(t)$ and then identify discrete spindle events using an empirical thresholding segmentation rule. We start with the classical chirp complex representation $x(t)=A(t)\exp(i\phi(t))$, where $A(t)$ is the instantaneous envelope that varies slowly in time, and $\phi(t)$ is the instantaneous phase capturing the fast oscillatory behavior. The envelope $A(t)$ can computed from the first step of the Empirical Mode Decomposition (EMD) [23, 21]. We define an isolated spindle pattern as an oscillatory events occurring between two consecutive local minima of the amplitude envelope, separated by a local maximum. Specifically, our segmentation procedure follows the following steps:
@@
    Noise reduction via smoothing: Estimate the upper $x_{\text{up}}(t)$ and lower $x_{\text{down}}(t)$ envelopes by applying locally weighted polynomial regression (LOESS) using a second-degree model for the local maxima and minima, respectively [35].
@@
Computational complexity.
Using the sampling frequency $F_{s}$ and the window duration $T$ (Algorithm 1, $T=30$ s), so that $N=TF_{s}$ samples are processed per window. Using FFT-based implementations, both the Hilbert transform and autocorrelation estimation scale as $\mathcal{O}(N\log N)$ per window. The –not found– loop (run for $I$ iterations until acceptance) performs (i) one OU simulation of length $N$ and (ii) event segmentation/statistics extraction on the window, each in $\mathcal{O}(N)$ time, followed by a distributional comparison of $S$ detected events (e.g., KS/Wasserstein after sorting) in $\mathcal{O}(S\log S)$ time with $S\leq N$. Overall, the per-window time complexity is thus
$\mathcal{O}\!\left(N\log N+I\,(N+S\log S)\right)\subseteq\mathcal{O}\!\left((1+I)\,N\log N\right),$ (16)
and the memory is $\mathcal{O}(N)$. Using the segmentation describe above, we extract the spindle duration, peak amplitude, and mean frequency.
@@
where “p.v.” is the Cauchy principal value. From the conjugate component $y_{k}(t)=\mathcal{H}\{x_{k}\}(t)$, we have $x_{a,k}(t)=x_{k}(t)+iy_{k}(t)=A_{k}(t)\exp(i\phi_{k}(t))$, where $A_{k}(t)$ is the instantaneous amplitude and $\phi_{k}(t)$ is the instantaneous phase of spindle $k$. The instantaneous frequency is given by $\omega_{k}(t)=\frac{\mathrm{d}}{\mathrm{d}t}\phi_{k}(t)$, which is a pointwise measure of oscillatory activity within the spindle [36]. The mean frequency of spindle $k$, segmented between time $t_{k}$ and $t_{k+1}$ is the average over its duration $T_{k}=t_{k+1}-t_{k}$:
@@ III-A General projection on two-dimensional OU-processes
We apply here the OU-decomposition to the EEG signal during general anesthesia, mostly dominated by two main frequency bands: $\delta$ (0.5-4 Hz) and $\alpha$ (8-14 Hz). These bands result from random spindle patterns (Fig. 3A-C).
Figure 3: Spectral decomposition of EEG in $\alpha$- and $\delta$- and residual components A. EEG and spectrogram over several seconds (scal bar 5s), showing $\alpha$- and $\delta$- spindles. B. $\alpha$- and $\delta$- filtered components. C. Power spectral decomposition (PSD). We fitted the PSD in intermediate frequency bands $[1.8,4]$ and $[6.5,8.1]$ Hz with $\log PSD=a\log(f)+b$. D. Statistical result of (a,b) coefficients for the two bands and Lorentzian parameters $\frac{A\sigma}{\pi}\frac{1}{\sigma^{2}+(f-f_{c})^{2}}$, computed over n=18 EEGs.
@@
to the PSD for each band and obtain the distributions shown in Fig. 3D. Note that the PSD is computed on a window large enough (here  5min). Finally, the scaling coefficient for the two decaying component is compued by fitting the PSD in intermediate frequency bands $[1.8,4]$ (resp. $[6.5,8.1]$)Hz with $\log PSD=a\log(f)+b$, $a=-.7\pm 0.5$ (resp. $a_{high}=-4.5\pm 0.49$). Note that this parameter estimation directly gives similar result compared to the three functions method introduced above.
To conclude we proposed here a decomposition of the EEG signal with OU-processes centered around a dominant frequency band. This approximation is valid on some restricted intervals, separated by scaling law in between.
@@ IV Estimating the OU-parameters based on spindle statistics
Figure 4: Parameter estimation of the stochastic OU- model to EEG spindles via grid search. (A) Parameter estimation pipeline. A grid search is performed over the parameter space of the two-dimensional Ornstein–Uhlenbeck process ($\lambda$, $\sigma$), while fixing the resonance frequency $f_{c}=\omega/2\pi$. (B) Grid search minimizing a loss function based on spindle statistics with (duration, amplitude) computed from segmented EEG recordings, this leads to a line of minimum with slope $\lambda/\sigma=0.082$. However, when considering the spindle statistics with (number, duration, and amplitude), the optimal parameters are $\lambda^{*}=1.84$, $\sigma^{*}=35.59$, and $f_{c}^{*}=9.63$ Hz. (C) EEG signal (top) and stochastic simulations of the OU process (bottom) using the optimal parameters. (D) Spindle segmentation applied to both real EEG and simulated signals. Events are identified and classified based on amplitude and duration thresholds. (E) Comparison of spindle statistics (number, amplitude, and duration) between EEG and simulations.
Using a square grid search, for a fixed ensemble $(\lambda,\sigma)$ with predefined boundary, the method consists in three steps: generating OU-realizations, segmenting their spindles to obtain the duration and amplitude distribution, and comparing them with the distribution of the empirical signal (Fig. 4A). The next point is obtained by starting again from $(\lambda+\delta\lambda,\sigma+\delta\sigma)$. To compare the two distribution with the empirical one, we use the Wasserstein distance [37] based on the sum of absolute differences between the two cumulative density functions
@@
The optimization procedure that allows to identify the optimal parameters is
@@
The results of such a procedure (Fig. 4B) depends on the parameter to be considered: when considering the duration and amplitude only in eq. 27, we do not find any single optimal parameter, due to a degeneracy (duration is proportional to amplitude as shown in Fig. 4B-Last raw. In that case, we obtain a line of minima where the ratio ${\sigma}/{\lambda}$ is constant (fig. 4B). However, when adding the number of spindles to the two other parameters, we remove the degeneracy and obtain a global minimum for the loss function $S_{\lambda,\sigma}$, associated with the optimal parameters (fig. 4B). Finally, we obtain a similar trend between the amplitude and duration distribution of the EEG signal vs. the optimal reconstructed OU-processes. To conclude such second approach allow to estimate OU-parameters based on the statistics of spindles.